OpenStudy (anonymous):

i need help with these two problems SUB FORMULA evaluate the integral problem one ∫0π(1+cos5t)2sin5tdt problem two ∫π/6π/2cotxcsc3xdx

OpenStudy (anonymous):

i know the D, DU

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} \] \[\int\limits_{\pi/6}^{2/\pi}\] to make it more clear

OpenStudy (anonymous):